OpenStudy (anonymous):

Solve by using the Quadratic Formula. Write both solutions separated by a comma: 2x^2-9x+9=0

OpenStudy (bahrom7893):

D = 81-72=9 x1 = (9-3)/4= 6/4 = 3/2 x2 = (9+3)/4 = 12/4 = 3